OpenStudy (anonymous):

i have to solve the polynomail equation by factoring. This is the problem 9x2+36x-189. (The 2 signifies "squared". I get to 9(x+7)(x-3) and I am lost. Someone help please :]

OpenStudy (amistre64):

what is the 'solution' you are needing?

OpenStudy (anonymous):

i must solve the polynomial equation.

OpenStudy (amistre64):

what does 'solve' mean in this context?

OpenStudy (anonymous):

as in for zero products

OpenStudy (amistre64):

ok.. then you are looking for the zeros, or roots of the equation; and not just the factored form right?

OpenStudy (amistre64):

9x2+36x-189 ; /9 9 (x^2 +4x -21) ; the 9 is useless to us in finding zeros so toss it x^2 +4x -21 ; factor it

OpenStudy (amistre64):

(x+7)(x-3) are good now we want to know when either one of them equals ZERO

OpenStudy (amistre64):

x+7 = 0; when x = ? x-4 = 0; when x = ?

OpenStudy (anonymous):

im sorry after reviewing my notes, it can aslo be solving a quadriatic equation by factoring. it must be factored out. A more simpler problem would be: \[n ^{2} +9n+14=0...n ^{2} turns into (n+7)(n+2).. therfore, n+2=0 \in order \to get n by itself subtract 2 from each side leaving you with n=-2.. again with n+7=0 \in ordcer \to get n alone subtract 7 from each side. n=-7.. The set is (-2,-7)\]

OpenStudy (anonymous):

makes no sense lol

OpenStudy (amistre64):

9(x+7)(x-3) is a good 'solution' then

OpenStudy (anonymous):

Ok I understand up to that Im just just confused as to where the 9 goes.

OpenStudy (amistre64):

9 = 0 never, so it plays no part in finding zeros; other than that, it just sits on the outside looking in

OpenStudy (anonymous):

No, 9(x+7)(x-3) is not the solution.. Ok Thank you for your help I may just review my lesson over again because I don;t beleive we would just toss the 9 out like that. it is still part of the problem and when checking my answer it does not match up if i just simply toss the 9. thank you anywy

OpenStudy (anonymous):

when you factor you leave it

OpenStudy (anonymous):

in fact u should always leave it

OpenStudy (amistre64):

the roots of the equation are the same whether you solve for 9=0 or not :)

OpenStudy (anonymous):

just in the last step

OpenStudy (anonymous):

you set all the factors in the product equal to zero

OpenStudy (anonymous):

ok like i said, Thank you anyway, amistre64! I aprrciate it! I think it just would be best for me to re-review my lesson considering i have doubts.

OpenStudy (anonymous):

with something like 9(x-3)(x+5) =0 you set 9=0 , or x-3 =0 or x+5=0

OpenStudy (amistre64):

(x+7)(x-3) has the same exact roots as 9(x+7)(x-3) which has the same exact roots as 174638485736738459858(x+7)(x-3)

OpenStudy (anonymous):

the first eqn is a false statement , you ignore it because its false, it gives no answer

OpenStudy (anonymous):

9=0 is not true for any x

OpenStudy (anonymous):

ok thank you elecengineer!!!!!!!!!!!!!!!!! that is exactly what i was confused about so you do set the 9 to 0! perfect! thank you so much!! :]

OpenStudy (anonymous):

it gives no extra solutions

OpenStudy (amistre64):

....... you can set 9 equal to 62 if you want and get the same results lol

OpenStudy (anonymous):

amister im sorry but you are severely confusing me furthger than i already am

OpenStudy (anonymous):

lol:P

OpenStudy (amistre64):

thats because im trying to teach you that you need to see the problem for what it is instead of just trying to apply blind methods.

OpenStudy (anonymous):

prob best to leave all the constant factors in , until the very last step , instead of just tossing them away. Even though, yes , you know what you are doing , it is a tad confusing to someone that doesnt really know what they are doing.

OpenStudy (amistre64):

set 9 = 0 is nonsense language; 9=9 and can not be "set" to 0; since 9=9 it is not a 'zero' of the function....ever :)

OpenStudy (anonymous):

amistre you are being a little to rude to me and making me feeling as if you are speaking down to me. please do not do that. I said ido not understand and rather than confusing myself further i will simply go over the lesson again.. i thanked you for your time and for your help that is all.

OpenStudy (anonymous):

or you could say "divide "

OpenStudy (anonymous):

ok thank you for your help and time

OpenStudy (shadowfiend):

Hokay now....

OpenStudy (anonymous):

but thats spliting hairs lols

OpenStudy (amistre64):

toss it, ignore it, put it in the corner, leave it out on the doorstep in the rain .... doesnt matter what you do with it, it serves no purpose in finding the roots of the function

OpenStudy (amistre64):

santani; trust me, if I were 'being rude' to you the whole world would no about it; but since your feeling upset about how you have read my responses; I will apologise.